The Directions Say "Solve The Following System Of Equations."
And The Problem Is.....
4x+6y=9
10x-18y=-5

Answer by checkley77(12844)

4x+6y=9 multiply by 3 & add.
10x-18y=-5
12x+18y=27
----------------
22x=22
x=22/22
x=1 ans.
4*1+6y=9
4+6y=9
6y=9-4
6y=5
y=5/6 ans.
Proof:
10*1-18*5/6=-5
10-90/6=-5
10-15=-5
-5=-5